On January 1, 2017, Corgan Company acquired 70 percent of the outstanding voting stock of Smashing, Inc., for a total of $700,000 in cash and other consideration. At the acquisition date, Smashing had common stock of $710,000, retained earnings of $260,000, and a noncontrolling interest fair value of $300,000. Corgan attributed the excess of fair value over Smashing's book value to various covenants with a 20-year remaining life. Corgan uses the equity method to account for its investment in Smashing.

During the next two years, Smashing reported the following:

Net Income Dividends Declared Inventory Purchases from Corgan
2017 $ 160,000 $ 36,000 $ 110,000
2018 140,000 46,000 130,000

Corgan sells inventory to Smashing using a 60 percent markup on cost. At the end of 2017 and 2018, 30 percent of the current year purchases remain in Smashing's inventory.

  1. Compute the equity method balance in Corgan's Investment in Smashing, Inc., account as of December 31, 2018.
  2. Prepare the worksheet adjustments for the December 31, 2018, consolidation of Corgan and Smashing.
Answer #1

Solution:

Investment account

Particulars Amounts($) Amounts($)
Consideration transferred 2017 $700,000
Smashing Net income(160,000 * 70%) 112,000
Covenants amortization(1,500 * 70%) -1,050
Closing inventory profit(100%) -12,375
Equity in Smashing Earnings(112,000-1,050-12,375) 98,575
2017 Dividend(36,000 * 70%) -25,200
Closing Investment Bal'17(700,000+98,575-25,200) 773,375
Smashing Net income 2018(140,000 * 70%) 98,000
Covenants amortization(1,500 * 70%) -1,050
Opening inventory profit 12,375
Closing inventory profit(100%) -14,625
Equity in Smashing earnings(98,000+12,375-1,050-14,625) 94,700
2018 Dividend (46,000*70%) -32,200
Closing investment Bal'18 835,875

Working Notes

Particulars Amount($)
Consideration transferred by Corgan 700,000
Non controlling interest fair value 300,000
Smashing acquisition date fair value(1+2) 1,000,000
Book value of subsidiary (710,000 + 260,000)    970,000
Excess fair over book value (3-4)      30,000
Excess assigned to covenants      30,000   
Useful life in years            20
Annual amortization(6/7) 1,500
2017 Ending Inventory Profit
Cost $110,000/1.6 = 68,750
Intra entity gross profit $110,000 - 68,750 = 41,250
Ending Inventory gross profit $41,250*30% = 12,375
2018 Ending Inventory Profit
Cost $ 130,000/1.6 = 81,250
Intra entity gross profit $130,000-81,250 = 48,750
Ending inventory gross profit $48,750*30% = 14,625